# Online Learning–Enhanced Lie Algebraic MPC for Robust Trajectory Tracking of Autonomous Surface Vehicles

### v1

@@ III-B Fourier-based Online Learning
III-C Offline Bi-level Feature Extraction
Inspired by the Fourier-based nature of wave and wind disturbances [lale2024falcon], the features we use are all Fourier-based. However, directly using a large number of random Fourier features for disturbance estimation leads to unnecessary computational overhead, while many of them fail to capture the dominant frequency components and spatial correlations of the residual term. To address this issue, we perform feature extraction to identify and retain components that contribute significantly to the residual representation. This reduces both the computational cost and the risk of overfitting while preserving model accuracy.
$\Phi_{f}(Z_{t})=\big[\sin(f_{j}\,z_{t,i}),\;\cos(f_{j}\,z_{t,i})\big]_{i=1,\dots,d;\,j=1,\dots,m}$ (23)
@@ III-C Offline Bi-level Feature Extraction
$\min_{\{f_{j}\}}\;\frac{1}{T}\sum_{t=1}^{\mathsf{T}}\ell\!\Big(h_{t},\;\hat{\mathbf{A}}^{\star}(f)\,\Phi_{f}(Z_{t})\Big).$ (24)
@@
$\hat{\mathbf{A}}^{\star}(f)\;=\;\arg\min_{\hat{\mathbf{A}}}\frac{1}{T}\sum_{t=1}^{\mathsf{T}}\ell\!\big(h_{t},\,\hat{\mathbf{A}}\,\Phi_{f}(Z_{t})\big)+\lambda\|\hat{\mathbf{A}}\|_{1}.$ (25)
with $\ell(\cdot,\cdot)$ denoting the L2 prediction loss. The inner problem optimizes the amplitude parameters $\hat{\mathbf{A}}$ with an $\ell_{1}$ regularization that promotes sparsity, while the outer problem adjusts the frequencies $\{f_{j}\}$ to minimize the overall prediction error.
The inner optimization is a convex problem with respect to $\hat{\mathbf{A}}$, which guarantees a unique global optimum for fixed frequencies. We solve the overall bilevel objective using alternating minimization: fix $\{f_{j}\}$ and update $\hat{\mathbf{A}}$ (inner loop), then fix $\hat{\mathbf{A}}$ and update $\{f_{j}\}$ (outer loop). This alternating scheme adaptively learns both the most informative frequencies and their corresponding amplitudes. Even in the presence of Gaussian noise in the disturbance data, the convexity of the inner layer ensures convergence to a stable global minimum. The implementation is detailed in Algorithm 2.
@@ IV-A2 Controller Tuning
We tune weighting matrices $Q$ and $R$ based on eigenvalues of ABKA-BK, avoiding complex conjugate pairs and ensuring real parts remain well below unity.
@@ IV-A3 Baselines
We compare the performance of our method with those baseline methods: (1) Nominal Lie-MPC: the Lie-MPC framework without online learning [jang2023convex]. (2) L1 Adaptive-MPC: the same MPC framework with an online L1-style disturbance–estimation module to calculate compensation based on residual dynamics as well. [1657243] (3) PID: A two-layer PID controller with an outer loop that computes the desired twist from pose errors, and an inner loop that tracks these references using two independent PID controllers:
$v_{x}^{\mathrm{ref}},\,\omega_{z}^{\mathrm{ref}}$ $=f_{\mathrm{outer}}(e_{x},e_{y},e_{\psi}),$ (28)
@@ IV-B2 Results
The PID controller exhibits the highest tracking error overall. Its performance degrades substantially with increasing wind speed, particularly in the zigzag trajectory, where the RMSE rises from approximately 2.9 at 0 m/s to nearly 8.8 at 3 m/s. Even in the simpler lawnmower trajectory, PID maintains errors above 2.1 across all wind conditions, showing limited adaptability to disturbances.
Similarly, the L1 adaptive MPC performs worse than nominal-MPC, possibly due to over-estimation under low disturbance. This pattern can also be observed in the lawnmower trajectory, where L1 performs poorly under low wind disturbance but outperforms nominal-MPC in 3 m/s wind.

### v2

@@ III-C Offline Bi-level Feature Extraction
1:Dataset $\{(Z_{t},h_{t})\}_{t=1}^{T}$; initial amplitudes $\hat{C}$; initial frequency set $\{f_{j}\}_{j=1}^{M}$; feature map $\Phi(\cdot)$; inner steps $n_{\mathrm{inner}}$; epochs NN; learning rate $\eta$; regularization $\lambda$
2:Output: Optimized Fourier feature map $\Phi(\cdot)$
3:for $\text{epoch}=1$ to NN do
4:  Freeze $\{f_{j}\}$; unfreeze $\hat{C} \triangleright$ Inner loop
5:  for $k=1$ to $n_{\mathrm{inner}}$ do
6:   for all $(Z_{t},h_{t})$ in dataset do
7:     $\hat{h}_{t}\leftarrow\hat{C}^{\top}\Phi(Z_{t})$
8:     $L_{\mathrm{inner}}\leftarrow\|\hat{h}_{t}-h_{t}\|_{1}+\lambda\|\hat{C}\|_{1}$ $\triangleright$ Compute loss
9:     $\hat{C}\leftarrow\hat{C}-\eta\nabla_{\hat{C}}L_{\mathrm{inner}}$ $\triangleright$ Update weights
10:   end for
11:  end for
12:  Freeze $\hat{C}$; unfreeze $\{f_{j}\}$ $\triangleright$ Outer loop
13:  for all $(Z_{t},h_{t})$ in dataset do
14:   $\hat{h}_{t}\leftarrow\hat{C}^{\top}\Phi(Z_{t})$
15:   $L_{\mathrm{outer}}\leftarrow\|\hat{h}_{t}-h_{t}\|_{1}$ $\triangleright$ Compute loss
16:   $\{f_{j}\}\leftarrow\{f_{j}\}-\eta\nabla_{f_{j}}L_{\mathrm{outer}}$ $\triangleright$ Update weights
17:  end for
18:end for
[…]
@@
$\min_{\{f_{j}\}}\;\frac{1}{T}\sum_{t=1}^{\mathsf{T}}\ell\!\Big(h_{t},\;\hat{C}(f)\,\Phi(Z_{t})\Big).$ (26)
@@
$\hat{C}(f)\;=\;\arg\min_{C}\frac{1}{T}\sum_{t=1}^{\mathsf{T}}\ell\!\big(h_{t},\,C\,\Phi(Z_{t})\big)+\lambda\|C\|_{1}.$ (27)
with $\ell(\cdot,\cdot)$ denoting the L2 prediction loss. The inner optimization loop optimizes the amplitude parameters $C$ with an $\ell_{1}$ regularization that promotes sparsity, while the outer optimization loop adjusts the frequencies $\{f_{j}\}$ to minimize the overall prediction error.
The inner optimization is a convex problem with respect to $C$, which guarantees a unique global optimum for fixed frequencies. We solve the overall bilevel objective using alternating minimization: fix $\{f_{j}\}$ and update $C$ (inner loop), then fix $C$ and update $\{f_{j}\}$ (outer loop). This alternating scheme adaptively learns both the most informative frequencies and their corresponding amplitudes. Even in the presence of Gaussian noise in the disturbance data, the convexity of the inner layer ensures convergence to a stable global minimum. The implementation is detailed in Algorithm 2.
@@ IV-A2 Controller Tuning
We tune weighting matrices $Q$ and $R$ based on eigenvalues of ABKA-BK, avoiding complex conjugate pairs and ensuring the norm remains well below unity.
@@ IV-A3 Baselines
We compare the performance of our method with those baseline methods: (1) Nominal Lie-MPC: the Lie-MPC framework without online learning [10]. (2) L1 Adaptive-MPC: the same MPC framework with an online L1-style disturbance–estimation module to calculate compensation based on residual dynamics as well. [4] (3) PID: A two-layer PID controller with an outer loop that computes the desired twist from pose errors, and an inner loop that tracks these references using two independent PID controllers:
$v_{x}^{\mathrm{ref}},\,\omega_{z}^{\mathrm{ref}}=f_{\mathrm{outer}}(e_{x},e_{y},e_{\psi}),\\$ (30)
@@ IV-B2 Results
The PID controller exhibits the highest tracking error overall. Its performance degrades substantially with increasing wind speed, particularly in the zigzag trajectory, where the RMSE rises from approximately 2.9 at 0 m/s to nearly 8.8 at 3 m/s. Even in the simpler lawnmower trajectory, PID maintains RMSE above 2.1 across all wind conditions, showing limited adaptability to disturbances. Similarly, the L1 adaptive MPC performs worse than nominal-MPC, possibly due to over-estimation under low disturbance. This pattern can also be observed in the lawnmower trajectory, where L1 performs poorly under low wind disturbance but outperforms nominal-MPC in 3 m/s wind.